OpenStudy (anonymous):

Boris's gas tank is 3/10 full. After he buys 5 gallons of gas, It is 1/2 full. How many gallons can Boris's tank hold?

jimthompson5910 (jim_thompson5910):

If x is the number of gallons the tank can hold, then \[\Large \frac{3}{10}x + 5 = \frac{1}{2}x\] is the equation you need to solve

OpenStudy (anonymous):

First, take the 1/2 and make it with the denominator of 10. So itd be 5/10. Correct?

OpenStudy (anonymous):

3/10 x +5 =1/2x right?